# 3DSS-Mamba: 3D-Spectral-Spatial Mamba for Hyperspectral Image Classification

### v1

@@ I Introduction
Comparatively, recent State Space Models (SSMs) turn to establishing long-distance dependency through state transitions, enjoying the promising attributes of linear computational complexity and scalability. As an effective alternative to the Transformer, Mamba [36] introduces the selective SSMs for 1-D sequence modeling along a specific orientation, which has demonstrated substantial potential in natural language processing (NLP) tasks [37]. To accommodate vision scenarios involving 2D-spatial awareness, Vim [38] and VMamba [39] extend the Mamba architecture by introducing a multi-directional scanning mechanism to achieve global contextual modeling, showcasing great efficiency and effectiveness in 2D visual tasks, such as object detection and semantic segmentation. Although Mamba architectures have demonstrated substantial potential in low-dimensional scenarios, the adaptability to high-dimensional HSI classification tasks involving 3D hyperspectral data requires further exploration.
@@
  • Extensive experiments are verified on public hyperspectral datasets. The results indicate the effectiveness and superiority of the proposed method.
@@ II-A Preliminaries
The developed Spectral-Spatial Token Generation module (SSTG) is constructed by a 3D convolution block and an embedding operation, which projects the HSI patch cube into spectral-spatial tokens. Taking the cropped pixel-wise patch cube $x\in{\mathbb{R}^{B\times B\times d}}$ as input, the tokenization process can be formulated as
@@ III-A1 Pavia University
The dataset was captured by the Reflective Optics System Imaging Spectrometer (ROSIS) over Pavia, Northern Italy. The imaging wavelength of the spectrometer ranges from 0.43 to 0.86 $\mu$m. After removing the noisy bands, the dataset encompasses 103 spectral bands and 610 $\times$ 340 pixels, with a spatial resolution of 1.3 m per pixel. There are a total of 42776 ground sample points, categorized into 9 types including Asphalt, Gravel, Trees, etc. The division details for training and testing sets are described in Table I.
@@
TABLE II: Ablation study on the accuracy metrics for different scanning routes in 3DSS.
Routes Pavia University
OA AA Kappa
1 Spectral-priority 98.34 97.12 97.81
2 Spatial-priority 99.18 98.14 98.91
3 Cross spectral-spatial 99.07 98.5 98.77
4 Cross spatial-spectral 99.32 98.74 99.14
5 Parallel spectral-spatial 99.34 98.95 99.12
TABLE III: Classification Accuracies of the Compared Methods in Terms of OA, AA, $\kappa$, and the Accuracies of Each class for the Pavia University Dataset. The Best Accuracies are Presented in Bold.
SVM 1D-CNN 2D-CNN 3D-CNN VIT SF HSI-BERT CASST
1 91.63 83.11 97.44 95.94 82.87 94.75 97.63 99.17 99.10
2 97.56 92.78 99.80 99.71 94.00 98.14 99.93 99.95 99.94
3 73.90 60.03 86.28 89.76 75.91 84.70 86.51 98.28 98.70
4 92.10 85.57 96.90 97.36 83.39 96.81 97.53 93.08 97.49
5 98.51 97.77 99.33 100.0 99.14 99.22 99.61 98.66 99.92
6 86.04 70.77 96.56 97.41 51.49 94.37 93.05 99.60 99.08
7 84.09 44.21 94.44 96.69 41.34 84.56 95.80 99.62 99.92
8 90.25 66.38 93.51 93.92 76.48 92.65 97.86 94.73 99.11
9 99.22 80.89 99.58 100.0 99.34 100.0 99.67 96.41 97.33
OA (%) 92.75 82.68 97.45 97.62 82.77 96.60 97.61 98.63 99.34
AA (%) 90.37 75.72 95.98 96.75 78.22 93.91 96.40 97.72 98.95
Kappa 90.35 76.92 96.61 96.84 76.77 94.17 96.82 98.19 99.12
Figure 3: Classification maps obtained by the compared methods on the Pavia University dataset. (a) Reference map. (b) SVM. (c) 1D-CNN. (d) 2D-CNN. (e) 3D-CNN. (f) VIT. (g) SF. (h) HSI-BERT. (i) CASST. (j) the proposed 3DSS-Mamba.
@@ III-B1 Evaluation Metrics
Following the state-of-the-art methods in HSI classification, overall classification accuracy (OA), average classification accuracy (AA), and kappa coefficient (Kappa) are adopted as the evaluation metrics. To guarantee fairness in comparison, all experiments are conducted under identical experimental conditions, and the reported results are derived as the averages from 5 consecutive experiments.
III-B2 Competitive Approaches
To validate the effectiveness of the proposed 3DSS-Mamba, three kinds of representative HSI classification approaches are selected for comparison, including conventional architectures (SVM [8]), CNN-based methods (1D-CNN [24], 2D-CNN [42], 3D-CNN [26]), and Transformer-based methods (VIT [31], SF [33], HSI-BERT [32], CASST [35]). The network architecture and experimental parameters are consistent with their respective original papers.
III-B3 Implementation Details
All the experiments are implemented on the PyTorch platform with one RTX 3090Ti GPU. The training epochs and batch size are set as 100 and 64, respectively. The Adam gradient descent optimizer with learning rate 0.001 is exploited for parameter optimization. The PCA dimension for reduction is set to 30.
@@ III-C1 Effectiveness of Different Scanning Routes in 3DSS
Acknowledging the impact of scanning dimension priority on modeling capability, this section explores the effectiveness of various constructed scanning routes, including Spectral-priority, Spatial-priority, Cross spectral-spatial, Cross spatial-spectral, and Parallel spectral-spatial. Table II illustrates the classification results in terms of accuracy metrics. As can be observed, all scanning routes achieve significant classification performance, demonstrating the superiority of 3DSS in modeling global spectral-spatial contextual relationships. Comparatively, the Parallel spectral-spatial route showcases the optimal result, surpassing the Spectral-priority by margins of 1.0%, 1.83%, and 1.31% for OA, AA, and Kappa, respectively. Consequently, the Parallel spectral-spatial route is selected for subsequent experiments.
III-D Experimental Comparison With Competitive Approaches
III-D1 Pavia University Dataset
@@
The classification experiments on the Pavia University dataset are conducted with 5% of the reference samples. Table III provides the quantitative comparison results with each competitive approach. As can be observed, the proposed 3DSS-Mamba yields the most competitive performance in comparison with other studied methods. Restricted by hand-crafted feature descriptors, conventional approaches exhibit limitations in handling HSI data with complex contextual semantics, resulting in unsatisfactory classification performance. Despite achieving encouraging results, CNN-based models struggle to establish long-range dependencies due to their limited receptive fields, failing to capture global spectral and spatial characteristics. In contrast to Transformer-based architectures, the proposed 3DSS-Mamba performs global spectral-spatial contextual modeling from the sequence modeling perspective, achieving relatively stable and superior performance across all categories. Compared to the suboptimal CASST, the quantitative improvements in terms of OA, AA, and Kappa up to 0.71%, 1.23%, and 0.93%, respectively.
@@
Additionally, the corresponding classification maps obtained by different approaches are visualized in Fig. 3. As can be observed, conventional methods such as SVM generally demonstrate noticeable noises owing to the limitations of feature extraction ability. The CNN-based and Transformer-based methods perform smoother and sharper, but still suffer from some detail misclassifications. In contrast, the proposed 3DSS-Mamba achieves the most consistent results with the ground truth, which presents the clearest category boundaries with minimal noise.

### v2

@@ I Introduction
Figure 1: The motivation of the proposed 3DSS-Mamba. Mamba modeling has demonstrated substantial potential in low-dimensional scenarios such as 1D language and 2D visual tasks, motivating its adaptability to high-dimensional HSI classification task.
@@
Comparatively, recent State Space Models (SSMs) establish long-distance dependency through state transitions, enjoying the promising attributes of linear computational complexity and scalability. As an effective alternative to the Transformer, Mamba [36] introduces the selective SSMs for 1-D sequence modeling along specific orientation, which demonstrates substantial potential in natural language processing (NLP) tasks [37]. To accommodate vision scenarios involving 2D-spatial awareness, Vim [38] and VMamba [39] extend the Mamba architecture by introducing a multi-directional scanning mechanism to achieve global contextual modeling, showcasing great efficiency and effectiveness in 2D visual tasks, such as object detection and semantic segmentation. Although Mamba architectures have demonstrated substantial potential in low-dimensional scenarios, the adaptability to high-dimensional HSI classification tasks involving 3D hyperspectral data requires further exploration, as depicted in Fig. 1.
@@
  • Extensive experiments are verified on three public hyperspectral datasets. The results indicate the effectiveness and superiority of the proposed method.
@@ II-A Preliminaries
Figure 4: The detail structure of Spectral-Spatial Token Generation module (SSTG).
@@
The developed Spectral-Spatial Token Generation module (SSTG) is constructed by a 3D convolution block and an embedding operation, which projects the HSI patch cube into spectral-spatial tokens. The detailed structure is shown in Fig. 4. Taking the cropped pixel-wise patch cube $x\in{\mathbb{R}^{B\times B\times d}}$ as input, the tokenization process can be formulated as
@@ III-A Datasets Description
To illustrate the classification capabilities of the proposed 3DSS-Mamba, three publicly available HSI databases are utilized for comprehensive evaluation, including Pavia University, Indian Pines, and Houston 2013. Detailed descriptions are provided below.
@@ III-A1 Pavia University
The dataset was collected by the Reflective Optics System Imaging Spectrometer (ROSIS) over Pavia, Northern Italy. The imaging wavelength of the spectrometer ranges from 0.43 to 0.86 $\mu$m. After removing the noisy bands, the dataset consists of 103 spectral bands and 610 $\times$ 340 pixels, with a spatial resolution of 1.3 m per pixel. There are totally 42776 ground sample points, categorized into 9 types including Asphalt, Gravel, etc. Table I provides the division details for training and testing sets.
III-A2 Indian Pines
The dataset was acquired by the Airborne/Visible Infrared Imaging Spectrometer (AVIRIS) imaging an Indian pine tree over Northwestern Indiana in 1992. The imaging wavelength of the spectrometer ranges from 0.4 to 2.5 $\mu$m. The dataset encompasses 200 spectral bands and 145 $\times$ 145 pixels after removing the water absorption channels, with a spatial resolution of 20 m per pixel. There are a total of 10249 ground object pixels, representing 16 distinct categories including Alfalfa, Corn-notill, etc. The detailed data division in the experiment is described in Table II.
III-A3 Houston 2013
The dataset was captured by the ITRES CASI-1500 sensor over the University of Houston campus and its surrounding areas, provided by the 2013 GRSS Data Fusion Contest. The image comprises 144 spectral bands ranging in wavelength from 0.38 to 1.05 $\mu$m, and consists of 340 $\times$ 1905 pixels with a spatial resolution of 2.5 m per pixel. There are 16373 sample pixels, covering 15 challenging land cover categories. The precise splitting of training and testing data is exhibited in Table III.
@@ III-A3 Houston 2013
TABLE II: Name and Number of Samples of Each Class on the Indian Pines Datasets with 10% labeled data.
Indian Pines dataset No. Class Name Train Test 1 Alfalfa 5 41 2 Corn-notill 143 1285 3 Corn-mintill 83 747 4 Corn 24 213 5 Grass-pasture-mowed 48 435 6 Grass-trees 73 657 7 Grass-pasture 3 25 8 Hay-windrowed 48 430 9 Oats 2 18 10 Soybean-notill 97 875 11 Soybean-mintill 245 2210 12 Soybean-clean 59 534 13 Wheat 20 185 14 Woods 126 1139 15 Buildings 39 347 16 Stone 6 84 Total 1024 9225
TABLE III: Name and Number of Samples of Each Class on the Houston 2013 Datasets with 10% labeled data.
Houston 2013 dataset No. Class Name Train Test 1 Healthy grass 125 1238 2 Stressed grass 125 1241 3 Synthetic grass 70 690 4 Trees 124 1231 5 Soil 124 1229 6 Water 33 321 7 Residential 127 1255 8 Commercial 124 1231 9 Road 125 1239 10 Highway 123 1214 11 Railway 123 1222 12 Parking Lot 1 123 1220 13 Parking Lot 2 47 464 14 Tennis Court 43 423 15 Running Track 66 653 Total 1502 14871
Figure 5: Sensitivity analysis for the proposed method with different sizes of input patches.
TABLE IV: Ablation study on the accuracy metrics for different scanning routes in 3DSS.
Routes Pavia University Indian Pines Houston 2013
OA(%) AA(%) Kappa OA(%) AA(%) Kappa OA(%) AA(%) Kappa
1 Spectral-priority 98.34 97.12 97.81 93.21 84.73 92.25 97.95 98.07 97.78
2 Spatial-priority 99.18 98.14 98.91 95.49 88.06 94.85 98.31 98.39 98.17
3 Cross spectral-spatial 99.07 98.5 98.77 96.16 89.41 95.61 98.5 98.57 98.38
4 Cross spatial-spectral 99.32 98.74 99.14 96.16 90.22 95.62 98.84 99.04 98.86
5 Parallel spectral-spatial 99.34 98.95 99.12 96.47 93.09 95.97 98.93 98.92 98.85
Figure 6: Sensitivity analysis for the proposed method with different scales of 3D_Conv kernel in SSTG in terms of OA, AA, $\kappa$.
Figure 7: Sensitivity analysis for the proposed method with different embedding dims and depths for 3DMB. (a) Pavia University. (b) Indian Pines. (c) Houston 2013.
@@ III-B1 Evaluation Metrics
Following the state-of-the-art HSI classification approaches, overall classification accuracy (OA), average classification accuracy (AA), and kappa coefficient (Kappa) are employed as the evaluation metrics. To guarantee fairness in comparison, all experiments are performed under identical experimental conditions, and the reported results are averaged over five consecutive experiments.
III-B2 Implementation Details
All the experiments are implemented on the PyTorch platform with one RTX 3090Ti GPU. The training epochs and batch size are set as 100 and 64, respectively. The Adam gradient descent optimizer with learning rate 0.001 is exploited for parameter optimization. The PCA dimension for reduction is set to 30. Following the default hyperparameters in VMamba [39], the state dimension and expansion ratio in the 3DSS mechanism are fixed at 16 and 2, respectively.
@@ III-C1 Effectiveness of Different Scanning Routes in 3DSS
Acknowledging the impact of scanning dimension priority on modeling capability, this section explores the effectiveness of constructed five scanning routes, including Spectral-priority, Spatial-priority, Cross spectral-spatial, Cross spatial-spectral, and Parallel spectral-spatial. Table IV illustrates the classification results in terms of accuracy metrics. As can be observed, all scanning routes achieve significant classification performance, demonstrating the superiority of 3DSS in modeling global spectral-spatial contextual relationships. Comparatively, the Spatial-priority scanning route demonstrates more competitive advantages than the spectrum-prioritized mechanism. The integration of spatial and spectral information further contributes to the enhancement of classification capability. Notably, the Parallel spectral-spatial route showcases the optimal performance across all three datasets, benefiting from both the spatial and spectral priorities with bidirectional modeling. Taking the Pavia University dataset as an example, the Parallel spectral-spatial route surpasses the basic Spectral-priority by margins of 1.0%, 1.83%, and 1.31% for OA, AA, and Kappa, respectively. As a result, the Parallel spectral-spatial route is selected for subsequent experiments.
TABLE V: Classification Accuracies of the Compared Methods in Terms of OA, AA, $\kappa$, and the Accuracies of Each CLasses for the Pavia University Dataset. The Best Accuracies are Presented in Bold.
SVM 1D-CNN 2D-CNN 3D-CNN VIT(Pixel) VIT(Patch) SF HSI-BERT
1 91.63 83.11 96.31 95.94 82.87 92.46 94.75 97.63 99.33
2 97.56 92.78 99.37 99.71 94.00 96.53 98.14 99.93 99.34
3 73.90 60.03 82.22 89.76 75.91 93.73 84.70 86.51 95.09
4 92.10 85.57 94.35 97.36 83.39 95.64 96.81 97.53 95.91
5 98.51 97.77 99.85 100.0 99.14 100.0 99.22 99.61 99.69
6 86.04 70.77 93.65 97.41 51.49 98.47 94.37 93.05 99.43
7 84.09 44.21 89.17 96.69 41.34 94.62 84.56 95.80 95.88
8 90.25 66.38 88.73 93.92 76.48 98.60 92.65 97.86 96.20
9 99.22 80.89 97.57 100.0 99.34 98.11 100.0 99.67 97.22
OA (%) 92.75 82.68 95.77 97.62 82.77 96.18 96.60 97.61 98.48
AA (%) 90.37 75.72 93.47 96.75 78.22 96.46 93.91 96.40 97.56
Kappa 90.35 76.92 94.37 96.84 76.77 94.97 94.17 96.82 97.98
Figure 8: Classification maps obtained by the compared methods on the Pavia University dataset. (a) Reference map. (b) SVM. (c) 1D-CNN. (d) 2D-CNN. (e) 3D-CNN. (f) VIT(Pixel). (g) VIT(Patch). (h) SF. (i) HSI-BERT. (j) The proposed 3DSS-Mamba.
III-D Parameter Analysis
In this section, a series of experiments are carried out to analyze and determine the optimal parameters for 3DSS-Mamba, including the input patch sizes, the 3D convolution scales in SSTG, and the embedding dims and depths for 3DMB.
III-D1 Different Input Patch Sizes
Fig. 5 depicts the classification performance across different input patch sizes, ranging from $9\times 9$ to $17\times 17$ with a growth interval of 2. As observed from the figures, the Indian Pines and Houston 2013 exhibit a similar variation tendency of consistently increasing and then decreasing, with the maximum peak at the identified point $13\times 13$. For the Pavia University dataset, accuracy diminishes as the input patch size increases to 11. Accordingly, a patch size of $13\times 13$ is leveraged for the Indian Pines and Houston 2013 datasets, and $11\times 1$ 11 is employed for the Pavia University dataset.
TABLE VI: Classification Accuracies of the Compared Methods in Terms of OA, AA, $\kappa$, and the Accuracies of Each CLasses for the Indian Pines Dataset. The Best Accuracies are Presented in Bold.
SVM 1D-CNN 2D-CNN 3D-CNN VIT(Pixel) VIT(Patch) SF HSI-BERT
1 20.73 15.22 76.09 95.65 0.00 70.73 100.0 48.78 78.05
2 73.07 64.43 92.16 88.31 36.65 59.22 78.37 78.44 90.27
3 62.99 50.48 87.23 84.94 0.94 60.51 89.29 80.58 93.04
4 50.70 28.69 64.98 87.76 4.23 90.14 81.22 51.17 93.43
5 92.64 76.60 96.48 96.48 20.46 64.14 89.89 91.24 97.70
6 94.90 85.89 98.77 96.16 94.22 97.72 97.41 96.19 97.26
7 76.00 32.14 32.14 78.57 0.00 100.0 84.00 20.00 96.00
8 96.63 86.40 100.0 100.0 99.53 97.91 100.0 98.83 99.30
9 33.33 15.00 30.00 70.00 0.00 0.00 27.78 0.00 38.89
10 68.69 60.08 91.87 92.80 33.26 85.71 94.97 78.37 96.80
11 85.16 69.61 95.93 91.20 88.96 92.81 93.44 92.85 98.91
12 64.89 56.32 93.42 88.03 4.12 90.07 81.65 62.47 91.39
13 97.03 89.27 97.56 94.15 82.70 98.38 87.03 96.19 95.68
14 96.49 79.45 98.81 99.05 98.24 97.01 94.82 94.46 99.65
15 55.33 49.48 81.35 89.38 11.53 99.71 96.83 87.89 94.52
16 93.93 40.86 97.85 100.0 91.67 100.0 95.24 65.06 84.52
OA (%) 79.82 67.13 93.34 92.17 57.35 84.55 90.67 85.45 95.82
AA (%) 72.03 56.25 83.41 90.78 41.65 81.50 87.00 71.41 90.83
Kappa 76.84 62.44 92.39 91.08 48.99 82.32 89.36 83.36 95.23
Figure 9: Classification maps obtained by the compared methods on the Indian Pines dataset. (a) Reference map. (b) SVM. (c) 1D-CNN. (d) 2D-CNN. (e) 3D-CNN. (f) VIT(Pixel). (g) VIT(Patch). (h) SF. (i) HSI-BERT. (j) The proposed 3DSS-Mamba.
III-D2 Different Scales of 3D Convolution Kernel in SSTG
The generated 3D spectral-spatial tokens are determined by the scale of 3D convolution kernels within the spectral-spatial token generation module. Fig. 6 illustrates the classification sensitivity achieved with distinct 3D kernels on the three datasets. It can be observed that appropriately increasing the scale of 3D_Conv kernels contributes to capturing richer spectral-spatial contextual information. Based on the results, the optimal kernel scale for all three datasets is established as 32 & [3, 5, 5].
III-D3 Different Embedding Dims and Depths for 3DMB
As the core of 3DSS-Mamba, the 3DMB module based on 3DSS scanning mechanism is iteratively stacked to achieve the extraction of global spectral-spatial semantic representations. To explore the optimal structure of 3DSS-Mamba for classification, mixed experiments are carried out by simultaneously adjusting the embedding dimensions in 3DSS and the stacked depth of 3DMB. Fig. 7 demonstrates the classification sensitivity on three datasets. The range of embedding dim is set to $\left\{{8,16,24,32,48}\right\}$, and the depth covers an interval of $\left\{{1,3,6,12,18}\right\}$. The optimal combination is highlighted with a green point. As observed across all three datasets, lower embedding dimensions can lead to performance degradation due to underfitting. Conversely, excessively high dimensions and deeper depths provide limited accuracy improvements but computational burdens. By trading off these metrics, the proposed 3DSS-Mamba is constructed as a lightweight structure, where the embedding dimension is determined as 32, and the stacked depth is selected as 1.
III-E Experimental Comparison With Competitive Approaches
To demonstrate the effectiveness of the proposed 3DSS-Mamba, three kinds of representative HSI classification architectures are selected for comprehensive comparison, including conventional methods (SVM [8]), CNN-based methods (1D-CNN [24], 2D-CNN [42], 3D-CNN [26]), and Transformer-based methods (VIT [31], SF [33], HSI-BERT [32]). The quantitative accuracies (OA (%), AA (%), and Kappa (%)) on the Pavia University, Indian Pines, and Houston 2013 datasets are summarized in Table V-VII, with the best results highlighted in bold. Corresponding visualization maps are provided in Fig 8-10.
TABLE VII: Classification Accuracies of the Compared Methods in Terms of OA, AA, $\kappa$, and the Accuracies of Each CLasses for the Houston 2013 Dataset. The Best Accuracies are Presented in Bold.
SVM 1D-CNN 2D-CNN 3D-CNN VIT(Pixel) VIT(Patch) SF HSI-BERT
1 98.53 90.25 98.40 97.04 93.34 91.56 99.29 92.26 99.02
2 98.05 95.14 99.20 98.48 99.29 94.69 98.05 83.15 99.56
3 98.88 100.0 100.0 100.0 100.0 100.0 100.0 99.20 100.0
4 97.81 93.89 98.31 98.95 87.41 98.48 96.43 93.47 97.68
5 98.57 93.40 99.28 99.44 99.37 92.75 100.0 99.73 99.91
6 92.64 99.38 92.62 100.0 97.26 100.0 95.89 78.76 98.29
7 91.98 80.84 94.01 93.38 85.28 97.11 96.06 91.06 97.55
8 91.96 70.26 89.63 94.29 67.50 91.52 91.70 86.68 96.16
9 85.09 67.65 83.87 93.61 70.63 87.67 94.59 89.60 96.27
10 93.16 67.64 94.62 96.33 79.89 89.04 97.46 89.03 99.55
11 87.01 69.64 90.53 96.84 58.63 93.88 90.56 95.31 98.20
12 87.25 62.94 89.38 95.46 51.44 87.84 99.28 88.63 98.38
13 32.35 49.04 77.83 94.24 6.64 96.68 56.40 93.12 96.68
14 98.91 97.20 99.30 98.60 79.48 99.74 98.44 97.92 99.48
15 85.08 99.09 100.0 99.85 98.65 99.66 99.66 99.66 100.0
OA (%) 91.74 81.06 93.93 96.77 79.27 93.63 95.46 91.67 98.37
AA (%) 90.03 82.42 93.80 97.10 78.32 94.70 94.25 91.84 98.44
Kappa 91.06 79.51 93.44 96.51 77.54 93.12 95.09 90.99 98.24
Figure 10: Classification maps obtained by the compared methods on the Houston 2013 dataset. (a) Reference map. (b) SVM. (c) 1D-CNN. (d) 2D-CNN. (e) 3D-CNN. (f) VIT(Pixel). (g) VIT(Patch). (h) SF. (i) HSI-BERT. (j) The proposed 3DSS-Mamba.
III-E1 Pavia University Dataset
@@
The classification experiments on the Pavia University dataset are conducted with 5% of the reference samples. Table V provides the quantitative comparison results with each competitive approach. As can be observed, the proposed 3DSS-Mamba yields the most competitive performance in comparison with other studied methods. Restricted by hand-crafted feature descriptors, conventional approaches exhibit limitations in handling HSI data with complex contextual semantics, resulting in unsatisfactory classification performance. Despite achieving encouraging results, CNN-based models struggle to establish long-range dependencies due to their limited receptive fields. In contrast to Transformer-based architectures, the proposed 3DSS-Mamba performs global spectral-spatial contextual modeling from the sequence modeling perspective, achieving relatively stable and superior performance. Compared to the suboptimal approach, the quantitative improvements in terms of OA, AA, and Kappa up to 0.86%, 0.81%, and 1.14%, respectively.
TABLE VIII: Computational parameter analysis of various comparison approaches in terms of parameters, Flops, and inference time on the PU dataset.
Metrics 1D-CNN 2D-CNN 3D-CNN VIT(Pixel) VIT(Patch) SF HSI-BERT 3DSS-Mamba
Params (M) 0.0083 0.0101 0.0073 0.1107 0.1107 0.1222 0.3006 0.0103
Flops (G) 0.0015 0.0125 0.080 0.0139 1.1902 0.3292 5.4410 0.8936
Inference time (s) 0.91 0.72 1.13 1.64 3.49 6.55 37.46 4.67
Figure 11: Classification performance of the compared methods with different numbers of labeled data on three datasets. (a) Pavia University. (b) Indian Pines. (c) Houston 2013.
@@
Additionally, the corresponding classification maps obtained by different approaches are visualized in Fig. 8. As can be observed, conventional architecture such as SVM generally brings noticeable noises, which can be attributed to the limited feature extraction capability of hand-crafted descriptors. Typical CNN-based and Transformer-based approaches such as 1D-CNN and VIT(Pixel) suffer from significant misclassifications, particularly evident in the bare soil category. In contrast, the proposed 3DSS-Mamba achieves the most consistent results with the ground truth, which presents the clearest category boundaries with minimal noise.
III-E2 Indian Pines Dataset
The experiments on the Indian Pines dataset are performed with 10% of the reference samples. The quantitative classification accuracies are reported in Table VI. Based on the results, the proposed 3DSS-Mamba achieves the highest recognition performance in comparison with other competitive approaches, and exhibits excellent improvements in several categories, such as Corn-notill, Corn-mintil, and Soybean-notill. Due to the lack of consideration of spatial information, the Transformer-based VIT(Pixel) method inevitably suffers from undesirable classification consequences, with the reduction compared to 3DSS-Mamba reaching 38.47% for OA, 49.18% for AA, and 46.24% for Kappa, respectively. To highlight the differences in classification results, Fig 9 further provides the visualization maps of various methods. Benefiting from the extraction of global spectral-spatial semantic information with the 3D sequential modeling mechanism, the proposed 3DSS-Mamba exhibits the smoothest and clearest classifications across most regions despite slight edge information confusion. These phenomena further reveal the effectiveness and preeminence of the proposed method.
III-E3 Houston 2013 Dataset
The experiments on the Houston 2013 dataset are executed with 10% of the labeled samples. As evident from the classification results in Table VII, the proposed 3DSS-Mamba consistently outperforms other techniques by substantial margins, demonstrating the highest quantities across all three metrics. In contrast to the suboptimal method, 3DSS-Mamba achieves significant improvements in OA, AA, and Kappa by 1.6%, 1.34%, and 1.73%, respectively. Regarding the visualization maps in Fig 10, 3DSS-Mamba provides the most precise prediction details, even though this scenario is predominantly distributed with discrete and local sample targets. These excellent improvements further verifies the potential application of sequence scanning model in HSI classification.
Figure 12: Computational parameter analysis of various comparison approaches on the PU dataset.
III-F Analysis of Computational Complexity
This section investigates the computational complexity of the proposed 3DSS-Mamba, focusing on model parameters, Flops, and inference time. Fig 12 illustrates the model parameter sizes of various comparison approaches on the Pavia University dataset. Attributed to the inherent linear sequential modeling mechanism, the proposed 3DSS-Mamba achieves the best classification performance with significantly fewer computational parameters. Although the CNN-based methods enjoy slight computational burdens, their capability of capturing long-range dependencies is constrained by their local receptive field, restricting further performance improvement. Besides, Transformer-based methods generally suffer from higher resource consumption due to the series of multi-head self-attention (MHSA) modules. While VIT(Patch) and SF methods provide competitive performance, their model parameters are almost 10 times that of 3DSS-Mamba. Table VIII further presents the detailed model parameters, Flops, and inference time. In summary, the proposed 3DSS-Mamba exhibits competitive advantages in balancing computational efficiency and classification effectiveness, highlighting significant potentiality and viability for HSI classification tasks.
III-G Robustness Assessment
To demonstrate the robustness of the proposed 3DSS-Mamba, extensive experiments are conducted considering various proportions of training samples. Specifically, the selected percentage for the Indian Pines and Houston 2013 datasets covers an interval of $\left\{{1.0\%,2.0\%,\ldots,10.0\%}\right\}$, and for the Pavia University is set to $\left\{{0.5\%,1.0\%,\ldots,5.0\%}\right\}$. Fig 11 displays the performance variations with different percentages on the four HSI datasets, with the proposed 3DSS-Mamba highlighted by red curves. There is a basically reasonable trend that the classification accuracy of 3DSS-Mamba steadily increases with the percentages of training samples, which exhibit substantial robustness. Furthermore, 3DSS-Mamba consistently outperforms other competitive methods across most percentages on the three testing datasets. These notable advantages further demonstrate the feasibility and superiority of the proposed method.